please check this for me? I have to factor completely.
49x^2-25
-25+79x^2=0
-25+25+49x^2=0+25
0+49x^2=0+25
49x^2=0+25
49x^2=25

Answer by stanbon(75887)

I have to factor completely.
49x^2-25 = (7x-5)(7x+5)
------------
-25+79x^2=0
79x^2-25=0
(sqrt(79)x -5)(sqrt(79)x+5)=0
---------------
-25+25+49x^2=0+25
Simplify to get 49x^2-25=0
(7x-5)(7x+5)=0
x=5/7 or x=-5/7
=============
0+49x^2=0+25
Simplify to get:
49x^2-25=0
same result
------------------
49x^2=0+25
49x^2=25
same result
=====================
Cheers,
Stan H.